# MeshAnything V2: Artist-Created Mesh Generation With Adjacent Mesh Tokenization

### v1

@@ Abstract
We introduce MeshAnything V2, an autoregressive transformer that generates Artist-Created Meshes (AM) aligned to given shapes. It can be integrated with various 3D asset production pipelines to achieve high-quality, highly controllable AM generation. MeshAnything V2 surpasses previous methods in both efficiency and performance using models of the same size. These improvements are due to our newly proposed mesh tokenization method: Adjacent Mesh Tokenization (AMT). Different from previous methods that represent each face with three vertices, AMT uses a single vertex whenever possible. Compared to previous methods, AMT requires about half the token sequence length to represent the same mesh in average. Furthermore, the token sequences from AMT are more compact and well-structured, fundamentally benefiting AM generation. Our extensive experiments show that AMT significantly improves the efficiency and performance of AM generation.
Figure 1: Equipped with the newly proposed Adjacent Mesh Tokenization (AMT), MeshAnything V2 significantly surpasses MeshAnything (Chen et al., 2024b) in both performance and efficiency. MeshAnything V2 generates Artist-Created Meshes (AM) up to $1600$ faces aligned with given shapes. Combined with various 3D asset production pipelines, it efficiently achieves high-quality, highly controllable AM generation.
@@ 1 Introduction
To address this issue, very recently, a line of work (Nash et al., 2020; Alliegro et al., 2023; Siddiqui et al., 2023; Weng et al., 2024; Chen et al., 2024a; b) has focused on automatically generating Artist-Created Meshes (AMs) to replace manual labor. Inspired by the success of large language models (LLMs), these approaches treat AMs as sequences of faces and learn to generate them with autoregressive transformers (Vaswani et al., 2017) in a manner similar to LLMs. Unlike methods that produce dense meshes in a reconstruction manner, these methods learn from the distribution of meshes created by human artists, thereby generating AMs that are efficient, beautiful, and can seamlessly replace manually created meshes.
@@
Although these methods have achieved some success, they still face significant challenges. One major limitation is that current methods (Nash et al., 2020; Alliegro et al., 2023; Siddiqui et al., 2023; Weng et al., 2024; Chen et al., 2024a; b) cannot generate meshes with a large number of faces. Specifically, the maximum number of faces that can be generated is currently limited to $800$. The primary reason for this limitation is the inefficiency of the current tokenization methods. These methods treat a mesh as a sequence of faces, where each face consists of three vertices, and each vertex typically requires three tokens to represent. Consequently, each mesh is tokenized into a sequence nine times the number of its faces, resulting in substantial computational and memory demands. Besides, the resulting token sequence is highly redundant, which harms sequence learning and reduces performance.
This work aims to address this issue. We introduce a novel mesh tokenization method named Adjacent Mesh Tokenization (AMT), which processes meshes into more compact and well-structured token sequences, thereby improving both efficiency and performance. AMT achieves this by representing each face with one vertex instead of three whenever possible. As shown in Fig. 1 and Alg. 1, during the tokenization process, after encoding a face, AMT finds and encodes its adjacent face, which shares an edge, requiring only one additional vertex to represent the adjacent face. When an adjacent face cannot be found, AMT adds a special token & to the sequence to mark this event and restarts from a face that has not been encoded yet. Ideally, because AMT uses only one vertex to represent a face, the sequence length can be reduced to nearly one-third.
To test the effectiveness of AMT, we conducted extensive experiments in the MeshAnything (Chen et al., 2024b) setting. Our experiments on Objaverse (Deitke et al., 2023) demonstrated that AMT can, on average, reduce the sequence length by half, thereby reducing the computational load and memory usage of the attention block by nearly four times. Moreover, the performance of our model also improved due to the compact and well-structured token sequence from AMT. Besides, AMT can also be applied to unconditional or other conditional mesh generation settings (Siddiqui et al., 2023), and its effectiveness is not affected by the use of VQ-VAE (Van Den Oord et al., 2017).
Additionally, we introduce MeshAnything V2, which is equipped with AMT. V2 increases the maximum face count from $800$ to $1600$. Thanks to AMT, the performance and efficiency of V2 both significantly outperform its previous version (Chen et al., 2024b). We open-source MeshAnything V2 to the community.
@@
    We introduce a novel mesh tokenization method, named Adjacent Mesh Tokenization (AMT). Compared to previous tokenization methods, AMT requires approximately half the token sequence length to represent the same mesh, thereby fundamentally reducing the computational burden of Artist-Created Mesh generation.
@@
    We equip MeshAnything with AMT and introduce MeshAnything V2. V2 doubles the maximum number of faces that can be generated while achieving significantly better accuracy and efficiency.
@@
    Extensive experiments show that AMT significantly improves the efficiency and performance of mesh generation.
2 Related Work
@@
Sort the vertices of $\mathcal{M}$ in ascending order by their coordinates;
Sort the faces of $\mathcal{M}$ in ascending order by their vertex indices;
Initialize an empty token sequence $Seq$;
Initialize a face list $UnvisitedFaces$ as the faces of $\mathcal{M}$;
Remove the first face from $UnvisitedFaces$;
Append the three vertices of the removed face to $Seq$;
while $UnvisitedFaces$ is not empty do
       if the last token in $Seq$ is ”&” then
             Remove the first face from $UnvisitedFaces$;
             Append the three vertices of the removed face to $Seq$;
      else
             $AdjacentVertices\leftarrow$ vertices adjacent to both the last two vertices in $Seq$;
             Filter out vertices from $AdjacentVertices$ that form faces with the last two vertices in $Seq$ that are not in $UnvisitedFaces$;
             if $AdjacentVertices$ is not empty then
                   Sort $AdjacentVertices$ in ascending order by their coordinates;
                   Append the first vertex in $AdjacentVertices$ to $Seq$;
            else
                   Append ”&” to $Seq$;
return $Seq$;
@@ 2.1 Artist-Created Mesh Generation
Diverging from previous works that produce dense meshes, recent works have focused on generating meshes created by human artists, i.e., Artist-Created Meshes (AMs) (Nash et al., 2020; Alliegro et al., 2023; Siddiqui et al., 2023; Chen et al., 2024a; Weng et al., 2024; Chen et al., 2024b). These methods process meshes into ordered face sequences and learn to generate this sequence. (Nash et al., 2020) first proposed using autoregressive transformers to sequentially generate vertices and faces. (Siddiqui et al., 2023) use VQ-VAE to learn a mesh vocabulary and then learn this vocabulary with a decoder-only transformer. (Alliegro et al., 2023) differ from other methods by using a discrete diffusion model to generate AMs instead of an autoregressive transformer. (Chen et al., 2024a) propose directly using the discretized coordinates of the vertex as the token index, bypassing the need for VQ-VAE as in (Siddiqui et al., 2023). (Weng et al., 2024) use pivot vertices as a coarse mesh representation and then generate the complete mesh tokens. (Chen et al., 2024b) generate AMs aligned with given shapes, which can be integrated with various 3D asset production methods to convert their results into AMs.
@@
In recent years, 3D generation has gradually become one of the mainstream research directions in the field of 3D research. This area focuses on generating diverse, high-quality 3D assets for the 3D industry. Generative Adversarial Networks (GANs) (Wu et al., 2016; Achlioptas et al., 2018; Goodfellow et al., 2020) produce synthetic 3D data by training a generator and a discriminator network to distinguish between generated and real data. Very recently, a new line of works (Hong et al., 2023; Liu et al., 2024; Shi et al., 2023; Li et al., 2023; Tang et al., 2024; Wang et al., 2024; Tochilkin et al., 2024; Wei et al., 2024; Xu et al., 2024) directly generate 3D assets in a feed-forward manner. (Hong et al., 2023) pioneer these methods and use a transformer to directly regress the parameters of 3D models given conditions. Besides, applying diffusion models (Ho et al., 2020) to directly generate 3D assets has also been widely researched (Zhou et al., 2021; Nichol et al., 2022; Alliegro et al., 2023; Lyu et al., 2023; Liu et al., 2023; Zhang et al., 2024). (Zhang et al., 2024) lead the SOTA of current 3D generation methods by first generating high-quality 3D shapes with DiT and then producing detailed textures with material diffusion models.
As mesh is a crucial component in 3D generation, Artist-Created Mesh Generation (Nash et al., 2020; Alliegro et al., 2023; Siddiqui et al., 2023; Chen et al., 2024a; Weng et al., 2024; Chen et al., 2024b) is closely related to 3D generation. However, it differs significantly from previous 3D generation methods as it mainly focuses on sequence learning, which is rarely seen in other 3D generation methods, to produce high-quality mesh topology.
@@ 3.1 Adjacent Mesh Tokenization
We first illustrate the tokenization methods used in previous methods (Nash et al., 2020; Alliegro et al., 2023; Siddiqui et al., 2023; Chen et al., 2024a; Weng et al., 2024; Chen et al., 2024b). Although there are slight differences in detail, the previous tokenization methods can be unified as follows: Given a mesh $\mathcal{M}$, vertices are first sorted in ascending order based on their z-y-x coordinates, where z represents the vertical axis. Next, faces are ordered by their lowest vertex index, then by the next lowest, and so on. The mesh is then viewed as an ordered sequence of faces:
@@
Substituting Equation equation 2 into Equation equation 1 gives:
$\mathcal{M}:=((v_{11},v_{12},v_{13}),(v_{21},v_{22},v_{23}),\ldots,(v_{N1},v_{N2},v_{N3}))=\text{Seq}_{V}$ (3)
@@
To resolve this issue, we propose Adjacent Mesh Tokenization (AMT) to obtain a more compact and well-structured $\text{Seq}_{V}$ than previous method. Our key observation is that the main redundancy of $\text{Seq}_{V}$ comes from representing each face with three vertices as in equation 2. This results in vertices that have already been visited appearing redundantly in $\text{Seq}_{V}$. Therefore, AMT aims to represent each face using only a single vertex whenever possible. As shown in Fig. 2 and Alg. 1, AMT efficiently encodes adjacent faces during tokenization, using only one additional vertex. When no adjacent face is available, as illustrated in the last step of Fig. 2, AMT inserts a special token & into the sequence to denote this event and restarts the process from a face that has not yet been encoded. To detokenize, simply reverse the tokenization algorithm as described in Alg. 1.
In the ideal case, where the special token ”&” is rarely used, AMT can reduce the length of $\text{Seq}_{V}$ obtained by previous methods to nearly one-third. Of course, in extreme cases, such as when each face in the mesh is completely disconnected from others, AMT performs worse than previous methods. However, since the datasets Deitke et al. (2023); Chang et al. (2015) used for AM generation are created by human artists, the meshes generally have well-structured topologies. Thus, the overall performance of AMT is significantly better than previous methods. As shown in Sec. 4.3, on the Objaverse test set, AMT can reduce the length of $\text{Seq}_{V}$ by half on average.
@@
Discussion on AMT and the use of VQ-VAE. After obtaining $\text{Seq}_{V}$, mesh generation methods then need to process it into a token sequence for sequence learning. (Siddiqui et al., 2023) propose to train a VQ-VAE (Van Den Oord et al., 2017) to achieve this. They take $\text{Seq}_{V}$ as input and learn a vocabulary of geometric embeddings with the VQ-VAE. After training the VQ-VAE, they then use the VQ-VAE’s quantized features as the input for the transformer (Vaswani et al., 2017). Very recently, (Chen et al., 2024a) proposed another method to process $\text{Seq}_{V}$ into a token sequence. They discard VQ-VAE and directly use the discretized coordinates of the vertex as the token index. It is important to emphasize that whether or not VQ-VAE is used does not affect AMT’s effectiveness. This is because AMT operates before the aforementioned methods. For example, in the case of using VQ-VAE, AMT first shortens the $\text{Seq}_{V}$ that represents $\mathcal{M}$, and the shortened $\text{Seq}_{V}$ is then quantized into an embedding sequence with the VQ-VAE.

### v2

@@ Abstract
Meshes are the de facto 3D representation in the industry but are labor-intensive to produce. Recently, a line of research has focused on autoregressively generating meshes. This approach processes meshes into a sequence composed of vertices and then generates them vertex by vertex, similar to how a language model generates text. These methods have achieved some success but still struggle to generate complex meshes. One primary reason for this limitation is their inefficient tokenization methods. To address this issue, we introduce MeshAnything V2, an advanced mesh generation model designed to create Artist-Created Meshes that align precisely with specified shapes. A key innovation behind MeshAnything V2 is our novel Adjacent Mesh Tokenization (AMT) method. Unlike traditional approaches that represent each face using three vertices, AMT optimizes this by employing a single vertex wherever feasible, effectively reducing the token sequence length by about half on average. This not only streamlines the tokenization process but also results in more compact and well-structured sequences, enhancing the efficiency of mesh generation. With these improvements, MeshAnything V2 effectively doubles the face limit compared to previous models, delivering superior performance without increasing computational costs. We will make our code and models publicly available.
Figure 1: Equipped with the newly proposed Adjacent Mesh Tokenization (AMT), MeshAnything V2 significantly surpasses MeshAnything [5] in both performance and efficiency. MeshAnything V2 generates Artist-Created Meshes (AM) up to $1600$ faces aligned with given shapes. Combined with various 3D asset production pipelines, it efficiently achieves high-quality, highly controllable AM generation.
* Corresponding authors.
@@ 1 Introduction
To address this issue, very recently, a line of work [15, 2, 19, 27, 4, 5] has focused on automatically generating Artist-Created Meshes (AMs) to replace manual labor. Inspired by the success of large language models (LLMs), these approaches treat AMs as sequences of faces and learn to generate them with autoregressive transformers [23] in a manner similar to LLMs. Unlike methods that produce dense meshes in a reconstruction manner, these methods learn from the distribution of meshes created by human artists, thereby generating AMs that are efficient, beautiful, and can seamlessly replace manually created meshes.
@@
Although these methods have achieved some success, they still face significant challenges. One major limitation is that current methods [15, 2, 19, 27, 4, 5] cannot generate meshes with a large number of faces. A primary reason for this constraint is the inadequacy of existing mesh tokenization methods. These methods treat a mesh as a sequence of faces, where each face consists of three vertices, and each vertex typically requires three tokens to represent. Consequently, each mesh is tokenized into a sequence nine times the number of its faces, resulting in substantial computational and memory demands. Besides, the resulting token sequence is highly redundant, which harms sequence learning and reduces performance.
Observing the above issues, we focus on the tokenization method of mesh generation in this work, aiming to improve the efficiency and quality of mesh generation. Extensive research in language models has demonstrated the importance of tokenization in sequence learning [29, 17, 10]. Moreover, unlike text, which inherently has a sequential structure, meshes are graph-based structures with 3D characteristics. For any given mesh, there are countless ways to represent it as a 1D token sequence, making the influence of tokenization methods even more pronounced for meshes. Therefore, we emphasize that research on mesh tokenization methods is of great importance.
The impact of mesh tokenization on autoregressive mesh generation can be considered in two main aspects. The first is efficiency: representing a mesh with shorter, more compact token sequences leads to reduced context length, thereby decreasing memory and computational complexity. The second aspect is the regularity of the token sequence. A shorter token sequence is not always better for mesh generation; the regularity and pattern consistency of the sequence are crucial for effective sequence learning. Effective mesh tokenization must balance both efficiency and regularity to achieve high-quality, efficient mesh generation.
Considering the above, we introduce MeshAnything V2, an advanced model for generating Artist-Created Meshes. This model incorporate several key improvements to boost both performance and efficiency. At the core of MeshAnything V2 is our novel Adjacent Mesh Tokenization (AMT) method. AMT optimizes the tokenization process by representing each face with a single vertex rather than the traditional three. As illustrated in Fig. 1 and Algo. 1, AMT encodes adjacent faces using just one additional vertex, largely reducing the sequence length. When an adjacent face cannot be identified, a special token ’&’ is introduced to indicate this interruption, allowing the model to resume from an unencoded face.
In previous mesh generation works, users were unable to control the number of faces generated by the model, often resulting in meshes that did not meet application requirements. To address this, we introduce a face count condition, allowing users to specify an approximate number of faces, ensuring that generated meshes align with desired specifications. Additionally, to enhance the robustness of AMT during inference, we incorporate Masking Invalid Predictions [15] to prevent the model from producing invalid tokens, such as generating another ’&’ token immediately after an ’&’ token.
Extensive experiments on the Objaverse [6] dataset demonstrate that AMT can halve the sequence length on average. This reduction translates to nearly a fourfold decrease in computational load and memory usage within the attention block. Furthermore, AMT remains effective across various mesh generation settings, including unconditional and conditional approaches [19], even when using VQ-VAE [22]. Finally, with the integration of AMT, MeshAnything V2 doubles the maximum number of faces generated compared to previous models, significantly boosting both performance and efficiency.
@@
    We introduce MeshAnything V2, a novel Artist-Created Mesh Generation model. V2 doubles the maximum number of faces that can be generated while achieving significantly better accuracy and efficiency.
@@
    The core of V2 is our newly proposed mesh tokenization method, Adjacent Mesh Tokenization (AMT). Compared to previous tokenization methods, AMT requires approximately half the token sequence length to represent the same mesh, thereby fundamentally reducing the computational burden of Artist-Created Mesh generation.
@@
    We conduct extensive experiments on various mesh tokenization methods to explore the essential properties required for effective mesh tokenization. Our experiments demonstrate that AMT significantly improves the efficiency and performance of mesh generation, and that the token sequence produced by mesh tokenization methods must balance compactness and regularity.
@@
1 Sort the vertices of $\mathcal{M}$ in ascending order by their coordinates;
2 Sort the faces of $\mathcal{M}$ in ascending order by their vertex indices;
3 Initialize an empty token sequence $Seq$;
4 Initialize a face list $UnvisitedFaces$ as the faces of $\mathcal{M}$;
5 Remove the first face from $UnvisitedFaces$;
6 Append the three vertices of the removed face to $Seq$;
7 Define a SpecialToken & to indicate break signal.
8while $UnvisitedFaces$ is not empty do
9       if the last token in $Seq$ is & then
10             Remove the first face from $UnvisitedFaces$;
11             Append the three vertices of the removed face to $Seq$;
12
13      else
14             $AdjacentVertices\leftarrow$ vertices adjacent to both the last two vertices in $Seq$;
15             Filter out vertices from $AdjacentVertices$ that form faces with the last two vertices in $Seq$ that are not in $UnvisitedFaces$;
16             if $AdjacentVertices$ is not empty then
17                   Sort $AdjacentVertices$ in ascending order by their coordinates;
18                   Append the first vertex in $AdjacentVertices$ to $Seq$;
19
20            else
21                   Append & to $Seq$;
22
23
24
25
26return $Seq$;
@@ 2.1 Artist-Created Mesh Generation
Diverging from previous works that produce dense meshes, recent works have focused on generating meshes created by human artists, i.e., Artist-Created Meshes (AMs) [15, 2, 19, 4, 27, 5]. These methods process meshes into ordered face sequences and learn to generate this sequence. [15] first proposed using autoregressive transformers to sequentially generate vertices and faces. [19] use VQ-VAE to learn a mesh vocabulary and then learn this vocabulary with a decoder-only transformer. [2] differ from other methods by using a discrete diffusion model to generate AMs instead of an autoregressive transformer. [4] propose directly using the discretized coordinates of the vertex as the token index, bypassing the need for VQ-VAE as in [19]. [27] use pivot vertices as a coarse mesh representation and then generate the complete mesh tokens. [5] generate AMs aligned with given shapes, which can be integrated with various 3D asset production methods to convert their results into AMs.
@@
In recent years, 3D generation has gradually become one of the mainstream research directions in the field of 3D research. This area focuses on generating diverse, high-quality 3D assets for the 3D industry. Generative Adversarial Networks (GANs) [28, 1, 7] produce synthetic 3D data by training a generator and a discriminator network to distinguish between generated and real data. Very recently, a new line of works [9, 12, 18, 11, 20, 25, 21, 26, 30] directly generate 3D assets in a feed-forward manner. [9] pioneer these methods and use a transformer to directly regress the parameters of 3D models given conditions. Besides, applying diffusion models [8] to directly generate 3D assets has also been widely researched [34, 16, 2, 14, 13, 31]. [31] lead the SOTA of current 3D generation methods by first generating high-quality 3D shapes with DiT and then producing detailed textures with material diffusion models.
As mesh is a crucial component in 3D generation, Artist-Created Mesh Generation [15, 2, 19, 4, 27, 5] is closely related to 3D generation. However, it differs significantly from previous 3D generation methods as it mainly focuses on sequence learning, which is rarely seen in other 3D generation methods, to produce high-quality mesh topology.
@@ 3 Method
In this work, we introduce a novel mesh tokenization method, named Adjacent Mesh Tokenization (AMT). Compared to previous tokenization methods, AMT reduces the token sequence length by approximately half to represent the same mesh, thereby significantly lowering the computational overhead of mesh generation. AMT is detailed in Section 3.1.
Next, in Section 3.2, we integrate AMT into MeshAnything [5] and introduce MeshAnything V2.
@@ 3.1 Adjacent Mesh Tokenization
We first illustrate the tokenization methods used in previous methods [15, 2, 19, 4, 27, 5]. Although there are slight differences in detail, the previous tokenization methods can be unified as follows: Given a mesh $\mathcal{M}$, vertices are first sorted in ascending order based on their z-y-x coordinates, where z represents the vertical axis. Next, faces are ordered by their lowest vertex index, then by the next lowest, and so on. The mesh is then viewed as an ordered sequence of faces:
@@
Substituting Equation (2) into Equation (1) gives:
$\mathcal{M}:=((v_{11},v_{12},v_{13}),\ldots,(v_{N1},v_{N2},v_{N3}))=\text{Seq}_{V}$ (3)
@@
To resolve this issue, we propose Adjacent Mesh Tokenization (AMT) to obtain a more compact and well-structured $\text{Seq}_{V}$ than previous method. Our key observation is that the main redundancy of $\text{Seq}_{V}$ comes from representing each face with three vertices as in (2). This results in vertices that have already been visited appearing redundantly in $\text{Seq}_{V}$. Therefore, AMT aims to represent each face using only a single vertex whenever possible. As shown in Fig. 2 and Algo. 1, AMT efficiently encodes adjacent faces during tokenization, using only one additional vertex. When no adjacent face is available, as illustrated in the last step of Fig. 2, AMT inserts a special token & into the sequence to denote this event and restarts the process from a face that has not yet been encoded. To detokenize, simply reverse the tokenization algorithm as described in Algo. 1.
In the ideal case, where the special token "&" is rarely used, AMT can reduce the length of $\text{Seq}_{V}$ obtained by previous methods to nearly one-third. Of course, in extreme cases, such as when each face in the mesh is completely disconnected from others, AMT performs worse than previous methods. However, since the datasets [6, 3] used for AM generation are created by human artists, the meshes generally have well-structured topologies. Thus, the overall performance of AMT is significantly better than previous methods. As shown in Sec. 4.3, on the Objaverse test set, AMT can reduce the length of $\text{Seq}_{V}$ by half on average.
Vertices Swap. Considering two faces $f_{1}$ and $f_{2}$:
$f_{1}=(v_{1},v_{2},v_{3}),f_{2}=(v_{1},v_{3},v_{4}).$
These two faces are connected by edge $(v_{1},v_{3})$. Assuming we first represent $f_{1}$ as $(v_{1},v_{2},v_{3})$, AMT will be interrupted because $f_{2}$ does not contain $v_{2}$, even though $f_{1}$ and $f_{2}$ are actually adjacent. To solve this issue, we introduce a special token \$ to swap vertices. When the \$ token appears in front of a vertex, it indicates that the next face will be formed by the first and the last vertex of the previous face, instead of the last two. For example, the token sequence $(v_{1},v_{2},v_{3},\$,v_{4})$ implies the mesh consists of two faces: $(v_{1},v_{2},v_{3})$ and $(v_{1},v_{3},v_{4})$, i.e. $f_{1}$ and $f_{2}$.
When swap is enabled, AMT can explore more adjacent faces when searching for the next face. Specifically, in line 13 of Algo. 1, $AdjacentVertices$ can be modified to additionally include vertices adjacent to the last and third-to-last vertices in $Seq$.
Although the swap operation introduces an additional special token, it reduces the number of interruptions and effectively shortens the token sequence. While adding a special token may potentially increase the difficulty of sequence learning, our experiments in Sec. 4.3 show that it has no noticeable impact.
@@
Discussion on AMT and the use of VQ-VAE. After obtaining $\text{Seq}_{V}$, mesh generation methods then need to process it into a token sequence for sequence learning. [19] propose to train a VQ-VAE [22] to achieve this. They take $\text{Seq}_{V}$ as input and learn a vocabulary of geometric embeddings with the VQ-VAE. After training the VQ-VAE, they then use the VQ-VAE’s quantized features as the input for the transformer [23]. Very recently, [4] proposed another method to process $\text{Seq}_{V}$ into a token sequence. They discard VQ-VAE and directly use the discretized coordinates of the vertex as the token index. It is important to emphasize that whether or not VQ-VAE is used does not affect AMT’s effectiveness. This is because AMT operates before the aforementioned methods. For example, in the case of using VQ-VAE, AMT first shortens the $\text{Seq}_{V}$ that represents $\mathcal{M}$, and the shortened $\text{Seq}_{V}$ is then quantized into an embedding sequence with the VQ-VAE.